Use the following information to answer question 3 and 4 Suppose that the current spot exchange rate is €0.80/$ and the bank quoted forward exchange rate is €0.7813/$. The one-year interest rate is 1.4% in the United States and 1.35% in France.

3. According to the Interest Rate Parity condition, what is the 1 year forward exchange rate?

a. €0.7817/$
b. €0.7809/$
c. €0.8004/$
d. €0.7996/$

4. What is your arbitrage strategy if you have $1,000,000 or €800,000?

a. $38,286.19
b. $23,757.6
c. €7,866.58
d. €18,561.8
Answer #1

As per IRPT, fair forward rate = spot rate*(1+Interest rate in France)/(1+Interest rate in US)

= 0.80(1+0.0135)/(1+0.014)

= Euro 0.7996/Dollar

i.e. d

4.Starting with $1,000,000

Convert into Euro at spot rate and get 1,000,000*0.8 = Euro 800,000

Invest and get 800,000(1+0.0135) = Euro 810,800

Convert into Dollar at forward rate and get 810,800/0.7813 = $1,037,757.58

Pay back Loan 1,000,000(1+0.014) = $1,014,000

Arbitrage Profit = $23,757.6

i.e. b